OpenStudy (anonymous):

Solve the equation. k+(2-5k)(6)=k+12

OpenStudy (dangerousjesse):

Solve for k: \(\large 6 (2-5 k)+k = k+12\) Distribute \(\textrm 6\) over \(\textrm{2-5 k.}\) \(\large 6 (2-5 k) = 12-30 k:\) \(\large k+12-30 k = k+12\) Combine like terms in \(\textrm{k-30 k+12.}\) \(\large k-30 k = -29 k:\) \(\large -29 k+12 = k+12\) Move terms with k to the left hand side. Subtract k from both sides: \(\large 12+(-29 k-k) = (k-k)+12\) Combine like terms in \(\textrm{-29 k-k.}\) \(\large -29 k-k = -30 k:\) \(\large -30 k+12 = (k-k)+12\) Look for two terms that sum to zero. \(\large k-k = 0:\) \(\large 12-30 k = 12\) Isolate terms with k to the left hand side. Subtract 12 from both sides: \(\large (12-12)-30 k = 12-12\) Look for two terms that sum to zero. \(\large 12-12 = 0:\) \(\large -30 k = 12-12\) Look for two terms that sum to zero. \(\large 12-12 = 0:\) \(\large -30 k = 0\) Divide both sides by a constant to simplify the equation. Divide both sides of \(\normalsize\textrm{ -30 k = 0 by -30:}\) \(\large\frac{-30 k}{-30} = \frac{0}{-30}\) Any nonzero number divided by itself is one. \(\large\frac{-30}{-30} = 1:\) \(\large k = \frac{0}{-30}\) Any number times zero is zero. \(\large\frac{0}{-30} = 0:\)

OpenStudy (dangerousjesse):

Which leaves you with...?

OpenStudy (anonymous):

Thank you so much